OpenStudy (anonymous):

Find the derivative of the function. g(t)= 7/ ln(t)

OpenStudy (amistre64):

looks simple enough :)

OpenStudy (amistre64):

7 Dt(1/ln(t))

OpenStudy (dumbcow):

use quotient rule: (f/g)' = f'g - g'f / g^2

OpenStudy (amistre64):

- 7 ------------ t ln^2(t)

OpenStudy (amistre64):

i hope that looks beter on your screen :)

OpenStudy (anonymous):

I tried this answer but it is incorrect. can u explain how u got this answer?

OpenStudy (amistre64):

7/ln(x) f' = (bt'-b't)/b^2 b = ln(x), b'=1/x , t=7, t'=0 ln(x)(0) - (7)1/x / ln^2(x) -7/x ln^2(x) might be able to simplify it more

OpenStudy (anonymous):

okay ill have to go ask for some more help b/c this is not a correct answer and i dont know y it is nt accepting my answer

OpenStudy (anonymous):

thnxs

OpenStudy (dumbcow):

No that is the correct answer, make sure you check that you're entering the (ln(x))^2 part correctly